OpenStudy (anonymous):

Express answer in exact form. Find the area of one segment formed by a square with sides of 6" inscribed in a circle. (Hint: use the ratio of 1:1:√2 to find the radius of the circle.) help me hero :(

OpenStudy (anonymous):

I explained most of this to you before...

OpenStudy (anonymous):

really idr my teacher had made me do it over again

OpenStudy (anonymous):

your teacher hates you

OpenStudy (anonymous):

maybe

OpenStudy (anonymous):

diameter of the circle is \[6\times \sqrt{2}\] radius of the circle is \[r=3\sqrt{2}\] area of the circle is \[\pi r^2=\pi(3\sqrt{2})^2=18\pi\] area of the square is \[6^2=36\] difference in the areas is \[18\pi-36\] and one fourth of that is \[\frac{18\pi-36}{4}\]

OpenStudy (anonymous):

since i am assuming that "one segment" means one of those 4 areas, that is your "exact" answer